OpenStudy (anonymous):

i dont get how to write out the equation of a line and find an equation of a line

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

=/

OpenStudy (anonymous):

Generally linear equations can be written in the form y=mx+c Where m is the gradient and c is the y-intercept The gradient can be found from \[(y _{1}-y _{2})/(x _{1}-x _{2})=m\] If you know two coordinates then c can be found by subbing in one pair of coordinates into y=mx+c

OpenStudy (anonymous):

Eliza is very correct, but you may be more familiar with the term slope. Gradient is just a certain calculus term that in this case refers to the slope of the line, m in the equation y=mx+b. Otherwise, follow Eliza. To find the slope, use his equation for the gradient if you are given two points. If you have a graph, you can use (rise)/(run), and count how many points upwards divided by how many points across to find the slope. To get the y-intercept, just find the place where the line intersects the y-axis. Plug those numbers into y=mx+b and you should have your answer.

OpenStudy (anonymous):

tysm

OpenStudy (anonymous):

best if you name a problem, then see how it is worked out